OpenStudy (anonymous):

(please help?) Given that ABCD is a rhombus, find x. A. –2 B. 2 C. 12 D. 24

OpenStudy (anonymous):



OpenStudy (anonymous):

8x-3=7x+11

OpenStudy (jdoe0001):

both sides are equal, so, yes, solve for "x"

OpenStudy (mathstudent55):

All sides of a rhombus are congruent, so do what @angie041397 did and solve the equation for x.

OpenStudy (anonymous):

X=14

OpenStudy (jdoe0001):

lol

OpenStudy (anonymous):

yes thats correct @ashleyvee

OpenStudy (jdoe0001):

@ashleyvee see above :)

OpenStudy (anonymous):

I still dont get the correct answer..

OpenStudy (anonymous):

?? am i doing something wrong

OpenStudy (anonymous):

8x-7x=11+3

OpenStudy (anonymous):

i keep getting x=14 but thats not one of my answers above..

OpenStudy (anonymous):

ohh..i know what i did wrong..

OpenStudy (anonymous):



OpenStudy (anonymous):

so its 8x-7x=13+11 ?

OpenStudy (anonymous):

x=24! Thank you angie :)

OpenStudy (mathstudent55):

yes